OpenStudy (anonymous):

When solving x2 - 12x = -13 by completing the square, what value is added to both sides of the equation? (3 points) Answer -6 6 36 144

OpenStudy (anonymous):

36 because by adding i can factor and get the square to be (x-6)^2

OpenStudy (mathstudent55):

a^2 + 2a + 1 = (a + 1)^2 x^2 + 6x + 9 = (x + 3)^2 x^2 - 18x + 81 = (x - 9)^2 The third coefficient (1, 9, 81 above) is the square of half of the middle coefficient (2/2)^2 = 1^2 = 1 (6/2)^2 = 3^2 = 9 (-18/2)^2 = (-9)^2 = 81